OpenStudy (anonymous):

a man can do certain task in 21 hrs, another can do the task in 28 hrs, and a boy can do the task in 48 hrs.find how long will it take to the task if all work together

OpenStudy (cwrw238):

work out 1/21 + 1/28 + 1/48 = 1/x where x is the time required

OpenStudy (fifciol):

in one hour first man will do 1/21 of his work second man 1/28 and a boy 1/48. so all of them will do 1/21 + 1/28 + 1/48 of the total work. Total work is 1/t so \[ \frac{ 1 }{ 21}+ \frac{ 1 }{ 28 }+ \frac{ 1 }{48}= \frac{ 1 }{t }\] solve for t

OpenStudy (anonymous):

..so what will be the answer ?

OpenStudy (cwrw238):

fifcol has explained it better

OpenStudy (cwrw238):

find the lcm of 21,28,48 and t and multiply each term by the lcm

OpenStudy (anonymous):

wat will be the lcm ?

OpenStudy (cwrw238):

first find prime factors of each number 21 = 3 * 7 28 = 2 * 2 * 7 48 = 2 * 2 * 2 * 2 * 3 the lcm is the lowest number into which all 3 numbers divide

OpenStudy (anonymous):

it means wat

OpenStudy (cwrw238):

lowest common multiple for example the lcm of 2 and 3 = 6 (2*3) the lcm of 6 and 2 = 6 ( not 12): 2 = 2 6 = 2 * 3 there are duplicate 2's so you just take one 2 lcm = 2*3 = 6

OpenStudy (fifciol):

or if you don't want to search for lcm you can simply take for lcm the product of denominators e.g. \[\frac{ 1 }{2 }+\frac{ 1 }{6 }=\frac{ 6 }{ 2*6}+\frac{ 2 }{2*6 }=\frac{ 6+2 }{12 }=\frac{ 8 }{12 }=\frac{ 2 }{3}\]

OpenStudy (cwrw238):

in case 0f 21, 28 and 48 first take 21 = 3* 7 28 = 2*2* 7 7 is common so lcm = 2*2*3*7 so we have 2 *2*3* 7 and 2*2*2*2*3 2*2 and 3 are common so lcm = 2 *2*2*2*3*7 = 16*21 = 336

OpenStudy (cwrw238):

ok so multiply each term by 336t 336t 336t 336t 336 --- + --- + ---- = 21 28 48 16t + 12t + 7t = 336 which is easy to solve

OpenStudy (cwrw238):

35t = 336 t = 336/35 grab your calculator and work this out